OpenStudy (anonymous):

Choose the missing step in the given solution to the inequality -x - 3 < 13 + 3x.

OpenStudy (anonymous):

Steps: 1) -x - 3 < 13 + 3x 2) 3) -4x < 16 4) X > -4

OpenStudy (anonymous):

Choices: A) -4x - 3 > 13 B) 2x - 3 < 18 C) -4x - 3 < 13 D) 2x - 3 > 18

OpenStudy (anonymous):

well the first step after you have the equation is to get the x's to one side. so you would have -4x-3<13 which is c

OpenStudy (anonymous):

Can you help me with this one ? Below are the steps to solve an equation: Step 1: |x - 4| + 1 = 8 Step 2: |x - 4| = 8 - 1 Step 3: |x - 4| = 7 Which of the following is a correct next step to solve the equation? A) x - 4 = -7 B) x + 4 = -7 C) -x - 4 = 7 D) x + 4 = 7

OpenStudy (anonymous):

Would it be D ?

OpenStudy (anonymous):

im not all that good at understanding the solve for x when it has absolute value bars sorry but I know someone that is @sodukogirl13

OpenStudy (anonymous):

Thanks anyways (:

OpenStudy (anonymous):

yep